According to the​ data, the mean quantitative score on a standardized test for female​ college-bound high school seniors was 500. The scores are approximately Normally distributed with a population standard deviation of 100. What percentage of the female​ college-bound high school seniors had scores above 637​? Answer this question by completing parts​ (a) through​ (g) below.

e. Use the Normal table to find the area to the left of the​ z-score that was obtained from a standardized test score of 637. The area to the left of the​ z-score is
Answer #1

e) Z = (X - mean) / sd = (637 - 500) / 100 = 1.37

P(X < 637) = P(Z < 1.37) = 0.9147 (ans)